Container A has 200 L of water, and is being filled at a rate of 6 litters per minute. Container B has 500 L of water and is being drained at 6 litters per minute. How many minutes, will it take for the two containers to have the same amout if water? and how do I figure that out?
Answer by lwsshak3(11628)

Container A has 200 L of water, and is being filled at a rate of 6 litters per minute. Container B has 500 L of water and is being drained at 6 litters per minute. How many minutes, will it take for the two containers to have the same amout if water? and how do I figure that out?
***
fill rate=1/6
drain rate=-1/6
let x=time it takes for the two containers to have the same amount if water
200+x/6=500-x/6
lcd:6
1200+x=3000-x
2x=1800
x=900
time it takes for the two containers to have the same amount if water=900 min
Check:
200+900/6=200+150=350 L
500-900/6=500-150=350 L